# A Survey of Cross-domain Graph Learning: Progress and Future Directions

### v1

@@ V-C Discussions
Mixture-oriented approaches mainly focus on simultaneously integrate structure and feature information across domains, enabling deep learning models to more comprehensively capture commonalities across domains. Nevertheless, the absence of a powerful feature alignment technique that effectively maps datasets from different domains into a shared feature space remains a significant limitation for feature-oriented approaches. As shown in Table I, earlier mixture-oriented approaches were only able to scale to limited cross-domain settings. Recently, some feature alignment techniques, such as large language models (LLMs), have been introduced, but they have only managed to achieve performance at a conditional scale. Indeed, a few newly proposed methods have demonstrated the potential to achieve open cross-domain scale, but their primary focus has been on exploring the feasibility of true open CDGL. For example, some approaches utilize domain-specific linear projection layers to align feature dimensions across domains, yet the resulting performance improvements have been limited.
@@ V-D Significance and Future Implications
Model Backbone Paradigm Scale Difficulty LLM Fine-tuning Prompting Domain Setting Task Code
Structure-oriented GCC[26] GIN Pre-training Open High - Social Networks, Academic Networks, etc Full-shot Node, Graph Link
PCRec[27] GIN Pre-training Limited High - Recommendation Full-shot Link -
APT[29] GIN Pre-training Open High - Citation Networks, Protein, etc Full-shot Node, Graph Link
GRADE[28] GCN Domain Adaptation Open Low - Citation Networks, Recommendation, etc Unsupervised Node, Link Link
FedStar[30] GCN, GIN Federated Learning Open Low - Bioinformatics, Social Networks, etc Full-shot Graph Link
ProCom[32] GCN Pre-training Conditional High - Social Networks, E-commerce, etc Few-shot Community Detection Link
@@
GraphControl[22] GIN Pre-training Open High - Social Networks, Academic Networks, etc Full/Few-shot Node, Graph Link
GA2E[23] GAE Pre-training Open Moderate - Citation Networks, Community Networks, etc Full-shot Node, Link, Graph -
BooG[31] GAT, GIN Pre-training Conditional High Sentence Transformer Citation Networks, Molecular, etc Zero/Few-shot Node, Link, Graph Link
UniAug[25] GCN, GIN Pre-training Open Moderate - Biological Networks, Chemical Networks, etc Full-shot Node, Link, Graph Link
OpenGraph[24] GT Pre-training Open High - Recommendation, Social Networks, etc Zero/Few-shot Node,Link Link
RiemannGFM[33] - Pre-training Open High - Citation Networks, Social Networks, etc Full/Few-shot Node,Link Link
Feature-oriented KTN[34] HGNN Domain Adaptation Limited Low - Academic Networks Zero-shot Node -
NaP[35] GCN Domain Generalization Limited High - Social Networks Full-shot Node -
CDFS-GAD[40] GraphSAGE Domain Adaptation Open Low - Social Networks Few-shot Anomaly Detection -
OFA[38] GCN, GAT, etc Pre-training Conditional High Sentence Transformer, etc Citation Networks, Molecular, etc Zero/Few-shot Node, Link, Graph link
ZeroG[41] GCN, GAT, etc Pre-training Conditional High Sentence Transformer, etc Citation Networks Zero-shot Node link
GraphAlign[39] GCN Pre-training Open High E5-small Citation Networks, Knowledge Graphs, etc Few-shot Node, Link link
RELIEF[37] GCN Pre-training Limited High - Citation Networks, Molecular, etc Few-shot Node, Graph link
GPF-plus[36] GIN Pre-training Limited Moderate - Citation Networks, Knowledge Graphs, etc Few-shot Node, Link link
DAGPrompT[42] GCN LoRA Limited High - Citation Networks, Molecular, etc Full/Few-shot Node,Graph Link
Mixture-oriented UDA-GCN[54] GCN Domain Adaptation Limited Low - Citation Networks Full-shot Node -
COMMANDER[56] GAT Domain Adaptation Limited Low - - Yelp, Amazon Unsupervised Anomaly Detection -
DA-GCN[92] DA-GCN Domain Adaptation Limited Low - E-commerce, Video Streaming Full-shot Node -
GCFL[57] GIN Federated Learning Open Low - Molecular, Social Networks Full-shot Graph -
METABrainC[60] GCN, GAT, etc Multi-task learning Limited High - PPMI, HIV, BP Few-shot Graph -
GAST[62] HGAT Domain Adaptation Limited Low GloVe, BERT Product Full-shot Sentiment classification -
CDGEncoder[58] GIN Domain Adaptation Open Low - Molecule, Bioinformatics, etc Few-shot Graph -
CrossHG-Meta[59] - Domain Generalization Limited High - Academic Networks Few-shot Node Link
CCDR[61] GAT Contrastive Learning Conditional High - Product Zero-shot Link Link
PTGB[43] GCN, GAT, etc Pre-training Limited High - Brain Network Full-shot Node, Graph Link
SDGA[44] GCN Domain Adaptation Limited Low - Citation Networks Semi-supervised Node -
ALCDR[69] LightGCN Domain Adaptation Conditional Low - Recommendation Full-shot Link -
DGASN[67] GAT Domain Adaptation Limited Low - Citation Networks Full-shot Link Link
CDTC[66] GIN Domain Adaptation Limited Low - Molecular Few-shot Graph -
ACT[65] GraphSAGE Domain Adaptation Limited Low - Social Networks Full/Few-shot Graph Link
DH-GAT[63] GAT Domain Adaptation Limited Low BERT Newswire, Social Media Full-shot Entity Recognition -
ALEX[64] GCN Domain Adaption Conditional Low - Citation Networks, Academic Networks Full-shot Node -
@@
PGPRec[70] GCN, GAT, etc Pre-training Conditional High - Recommendation Full-shot Link -
TAPE[45] GCN, etc Pre-training Conditional High - Citation Full-shot Node -
CrossLink[71] DyGFomer Domain Generalization Open High - Social Networks, Question Answered, etc Full-shot Link -
GCOPE[21] GCN, FAGCN Pre-training Open High - Citation Networks, Social Networks Few-shot Node, Graph Link
Uni-GLM[73] GCN, etc Pre-training Conditional Moderate BERT Citation Networks, E-commerce Networks Semi-supervised, Few-shot Node, Link Link
SOGA[46] GCN Domain Adaption Limited Low - Citation Networks Unsupervised Node -
STGP[68] STGNN Pre-training Open High - METR-LA, etc Few-shot Node link
GSPT[47] Transformer Pre-training Conditional Moderate SentenceBERT Citation Networks Few-shot Node, Link Link
ARC[48] - Pre-training Conditional High - Citation Networks, Social Networks, etc Few-shot graph anomaly detection link
UniGraph[49] GAT Pre-training Conditional High Llama, vicuna-7B, etc Citation, Molecular, etc Zero/Few-shot Node, Edge, Graph Link
TEA-GLM[79] GraphSAGE Pre-training Conditional High Vicuna-7B Citation network, e-commerce Zero-shot Node, Link link
GOFA[80] Transformer Pre-training Conditional High Llama2, Mistral, etc Citation and product Networks Zero-shot Node, Link Link
HiGPT[81] HGNN Pre-training Conditional High Sentence-BERT Movie, Academic Networks Few-shot/Zero-shot Node Link
MDGPT[72] GIN Pre-training Open High - Citation, E-commerce, etc Few-shot Node, Graph -
AnyGraph[50] GCN Pre-training Open High - E-commerce, Academic Networks, etc Zero-shot Node, Link, Graph Link
GIMLET[74] Transformer Pre-training Limited High T5 Molecular Few-shot/Zero-shot Graph Link
GraphTranslator[75] GraphSAGE Pre-traing Conditional High ChatGLM2-6B E-commerce, Academic Few-shot/Zero-shot Node Link
GraphGPT[76] Transformer, etc Instruction tuning Conditional High Vicuna-7B-v1.1/1.5 Citation Networks Zero-shot Node, Link Link
GITA[77] GCN, GAT Instruction tuning Open High GPT-4V, LLaVA-7B/13B Collaboration Networks, Citation, etc Zero-shot/Full-shot Node, Link Link
InstructGraph[78] - Instruction tuning Conditional High LLaMA2 GSM8K, NLGraph, etc Full/Few-shot Node, Link, Reasoning, etc Link
GraphCLIP[82] GCN, GAT Pre-traning Conditional High Qwen2-72B Academic,E-commerce,etc Few/Zero-shot Node, Link, Graph Link
GraphWiz[83] - Instruction tuning Limited High LLaMA2-7B/13B, Mistral-7B GraphInstruct, NLGraph, etc Zero-shot/Few-shot Reasoning Link
GraphLoRA[51] GAT LoRA Open High - e-commerce, Social Networks, etc Full/Few-shot Node Link
MDGFM[53] GCN Pre-training Open High - Citation Networks, Social Networks, etc Few-shot Node -
@@
SAMGPT[52] GCN Pre-training Open High - Citation Networks, Social Networks, etc Few-shot Node,Graph Link
@@
TABLE I: A summary of models that leverage cross-domain knowledge to enhance graph-related tasks in the literature, ordered by their release time. Paradigm denotes the core technique used to achieve Cross-Domain Graph Learning. Domain indicates the domains across which knowledge is transferred. Scale and Difficulty refer to the cross-domain scales and the level of difficulty. Fine-tuning denotes whether it is necessary to fine-tune the parameters of Graph Neural Networks (GNNs). Prompting indicates the use of graph- or text-formatted prompts. Acronyms in Task: Node refers to node-level tasks; Link refers to link-level tasks; Graph refers to graph-level tasks; Reasoning refers to Graph Reasoning.
VI Future Directions

### v2

@@ V-C Discussions
Mixture-oriented approaches mainly focus on simultaneously integrating structure and feature information across domains, enabling deep learning models to more comprehensively capture commonalities across domains. Nevertheless, the absence of a powerful feature alignment technique that effectively maps datasets from different domains into a shared feature space remains a significant limitation for feature-oriented approaches. As shown in Table I, earlier mixture-oriented approaches were only able to scale to limited cross-domain settings. Recently, some feature alignment techniques, such as large language models (LLMs), have been introduced, but they have only managed to achieve performance at a conditional scale. Indeed, a few newly proposed methods have demonstrated the potential to achieve open cross-domain scale, but their primary focus has been on exploring the feasibility of true open CDGL. For example, some approaches utilize domain-specific linear projection layers to align feature dimensions across domains, yet the resulting performance improvements have been limited.
@@ V-D Significance and Future Implications
VI Open Questions and Insights
Each of the three paradigms reflects a distinct philosophy on how transferable knowledge can be shared and generalized across heterogeneous graph domains. Building upon the systematic review of existing works, we further summarize shared insights and emerging challenges. Specifically, this section discusses three fundamental questions central to CDGL research: (1) What types of graph structural information (e.g., topology regularities) and feature representations are transferable across domains? This concerns identifying the domain-invariant components that generalize beyond domain-specific graph construction rules. (2) What is the nature of the knowledge being transferred across domains? Beyond explicit structures or features, transferable knowledge may manifest as higher-level statistical priors, semantic abstractions, or topological invariances that implicitly support domain generalization. (3) Can we truly achieve open cross-domain transfer across graphs that are not originally designed in graph form? This raises a critical question about whether a general-purpose graph foundation model can effectively bridge domains constructed from fundamentally different modalities, such as those derived from NLP or CV.
Recent advances across the three paradigms collectively indicate that both structural and feature information contain domain-invariant elements for cross-domain transfer. From the structural perspective, shared connectivity patterns such as local motifs, hierarchical topologies, or community structures tend to preserve universal relational principles, making structural transfer inherently easier even without domain-specific semantics. For example, when ignoring domain features, two graphs can be directly compared through kernel-based similarity measures [nikolentzos2021kernelForSharedStructure1], which standardize heterogeneous structures into comparable forms. This structural normalization allows models pretrained on source domains to more readily identify shared topological patterns in target domains.
While structural transfer captures domain-invariant topology, the unified transfer of structural and feature information is far more critical in real-world scenarios, as features assign concrete semantic meanings to different topological structures. This enables models to derive more fine-grained node or graph representations that are meaningful across diverse domains. However, achieving such unified transfer is highly non-trivial, which explains why numerous studies focus on learning how to align structural and feature spaces effectively [yuanmuchUnifiedImportance1]. In particular, textual or tokenized features become more transferable when combined with LLMs, as they naturally reside within a shared embedding space where semantic correlations can guide structural reasoning [qi2025bridgeglmLanguageTransfer1]. Through such representations, models can learn how specific textual semantics correspond to particular topological patterns and how structural configurations evolve as feature meanings shift. In essence, the transferable knowledge in CDGL arises from the latent correspondence between topology and semantics, reflecting how semantic variations reshape structural patterns. Achieving robust CDGL thus requires models capable of discovering the underlying design principles of graphs and harmonizing structural and semantic signals into invariant, domain-agnostic representations.
Beyond the above two questions, a more fundamental challenge lies in determining whether true open cross-domain transfer is achievable across graph data that are not originally designed in graph form. Existing open CDGL methods [zhao2024GCOPE] have primarily investigated transferring knowledge among various graph domains (e.g., homogeneous and heterogeneous graphs) that inherently follow graph construction principles. However, it remains unclear whether graphs derived from other modalities—such as those constructed from pure NLP corpora or CV datasets—can serve as compatible domains for transfer. Nevertheless, this challenge also highlights a potential opportunity: even for such secondarily designed graphs, the underlying design principles and the alignment between structural and semantic signals may still be transferable. For instance, text-attributed graphs are originally constructed from natural language data but have demonstrated transferability across textual domains [liu2024OFA, li2024zerog]. Similarly, visual graphs derived from image descriptions or region relationships can potentially leverage shared semantic representations through language grounding [wang2025multimodalgraphCVCDGL1]. These insights suggest that the boundary of graph domains may be extendable through cross-modal representations, offering a promising direction toward realizing genuine graph foundation models.
@@ VI Open Questions and Insights
Model GNN Backbone Paradigm Scale Difficulty LLM Fine-tuning Prompting Domain Setting Task Code
Structure-oriented
@@
GCC[qiu2020gcc] GIN Pre-training Open High - Social/Academic Networks, etc. Full-shot Node, Graph Link
@@
PCRec[wang2021pcrec] GIN Pre-training Limited High - Recommendation Networks Full-shot Link -
@@
APT[xu2023APT] GIN Pre-training Open High - Citation/Protein Networks, etc. Full-shot Node, Graph Link
GRADE[wu2023GRADE] GCN Domain Adaptation Open Low - Citation /Recommendation Networks, etc. Unsupervised Node, Link Link
FedStar[tan2023FedStar] GCN, GIN Federated Learning Open Low - Biological/Social Networks, etc. Full-shot Graph Link
ProCom[wu2024procom] GCN Pre-training Conditional High - Social/E-commerce, etc. Few-shot Community Detection Link
GraphControl[zhu2024graphcontrol] GIN Pre-training Open High - Social/Networks, etc. Full/Few-shot Node, Graph Link
GA2E[hu2024GA2E] GAE Pre-training Open Moderate - Citation/Community Networks, etc. Full-shot Node, Link, Graph -
BooG[cheng2024BooG] GAT, GIN Pre-training Conditional High Sentence Transformer Citation/Molecular, etc. Zero/Few-shot Node, Link, Graph Link
UniAug[tang2024UniAug] GCN, GIN Pre-training Open Moderate - Biological/Chemical Networks, etc. Full-shot Node, Link, Graph Link
GFT[wang2024gft] GraphSAGE Pre-training Open High - Social/Molecular Networks, etc. Few-shot Node, Link, Graph Link
OpenGraph[xia2024opengraph] GT Pre-training Open High - Recommendation/Social Networks, etc. Zero/Few-shot Node, Link Link
RiemannGFM[sun2025RiemannGFM] - Pre-training Open High - Citation/Social Networks, etc. Full/Few-shot Node, Link Link
GFSE[chen2025gfse] GT Pre-training Open High - Citation/Molecular Networks, etc. Few-shot Node, Graph Link
Feature-oriented
KTN[yoon2022KTN] HGNN Domain Adaptation Limited Low - Academic Networks Zero-shot Node -
NaP[wang2024NAP] GCN Domain Generalization Limited High - Social Networks Full-shot Node -
CDFS-GAD[chen2024CDFSGAD] GraphSAGE Domain Adaptation Open Low - Social Networks Few-shot Graph Anomaly Detection -
OFA[liu2024OFA] GCN, GAT, etc. Pre-training Conditional High Sentence Transformer Citation/Molecular Networks, etc. Zero/Few-shot Node, Link, Graph link
ZeroG[li2024zerog] GCN, GAT, etc. Pre-training Conditional High Sentence Transformer Citation Networks Zero-shot Node link
GraphAlign[hou2024graphalign] GCN Pre-training Open High E5-small Citation Networks, Knowledge Graphs Few-shot Node, Link link
RELIEF[zhu2024relief] GCN Pre-training Limited High - Citation Networks, Molecular, etc. Few-shot Node, Graph link
GPF-plus[fang2024gpfplus] GIN Pre-training Limited Moderate - Citation Networks, Knowledge Graphs, etc. Few-shot Node, Link link
DAGPrompT[chen2025DAGPrompT] GCN PEFT Limited High - Citation/Molecular Networks, etc. Full/Few-shot Node,Graph Link
UNPrompt[niu2024UNPrompt] GNN Pre-training Limited High - Social Networks, etc. Zero-shot Node Link
GADT3[pirhayatifard2025crossdomainGADT3] GraphSage Domain Adaptation Open High - Social Networks, etc. Full-shot Graph Anomaly Detection Link
Mixture-oriented
UDA-GCN[wu2020UDAGCN] GCN Domain Adaptation Limited Low - Citation Networks Full-shot Node -
COMMANDER[ding2021COMMANDER] GAT Domain Adaptation Limited Low - - Social Networks Unsupervised Graph Anomaly Detection -
DA-GCN[Guo2021DAGCN] GCN Domain Adaptation Limited Low - Recommendation Networks Full-shot Node -
GCFL[xie2021gcfl] GIN Federated Learning Open Low - Molecular/Social Networks Full-shot Graph -
METABrainC[yang2022METABrainC] GCN, GAT, etc. Multi-task learning Limited High - Brain/Molecular Networks, etc. Few-shot Graph -
GAST[zhang2022GAST] HGAT Domain Adaptation Limited Low GloVe, BERT Review Networks Full-shot Sentiment Classification -
CDGEncoder[hassani2022CDGEncoder] GIN Domain Adaptation Open Low - Molecular/Biological Networks, etc. Few-shot Graph -
CrossHG-Meta[zhang2022CrossHGMeta] HGNN Domain Generalization Limited High - Academic Networks Few-shot Node Link
CCDR[xie2022CCDR] GAT Contrastive Learning Conditional High - Recommendation Networks Zero-shot Link Link
PTGB[yang2023ptgb] GCN, GAT, etc. Pre-training Limited High - Brain Networks Full-shot Node, Graph Link
SDGA[qiao2023SDGA] GCN Domain Adaptation Limited Low - Citation Networks Semi-supervised Node -
ALCDR[zhao2023ALCDR] LightGCN Domain Adaptation Conditional Low - Recommendation Networks Full-shot Link -
DGASN[shen2023DGASN] GAT Domain Adaptation Limited Low - Citation Networks Full-shot Link Link
CDTC[zhang2023CDTC] GIN Domain Adaptation Limited Low - Molecular Networks Few-shot Graph -
ACT[wang2023ACT] GraphSAGE Domain Adaptation Limited Low - Social Networks Full/Few-shot Graph Link
DH-GAT[xu2023DHGAT] GAT Domain Adaptation Limited Low BERT News/Social Media Graphs Full-shot Entity Recognition -
ALEX[yuan2023alex] GCN Domain Adaptation Conditional Low - Citation/Academic Networks Full-shot Node -
PGPRec[yi2023PGPRec] GCN, GAT, etc. Pre-training Conditional High - Recommendation Networks Full-shot Link -
CrossLink[huang2024CrossLink] DyGFomer Domain Generalization Open High - Social Networks Full-shot Link -
GCOPE[zhao2024GCOPE] GCN, FAGCN Pre-training Open High - Citation/Social Networks Few-shot Node, Graph Link
Uni-GLM[fang2024uniglm] GCN Pre-training Conditional Moderate BERT Citation/E-commerce Networks Semi-supervised, Few-shot Node, Link Link
SOGA[mao2024SOGA] GCN Domain Adaptation Limited Low - Citation Networks Unsupervised Node -
STGP[hu2024stgp] Graph Transformer Pre-training Open High - Traffic Networks Few-shot Node link
GSPT[song2024GSPT] - Pre-training Conditional Moderate SentenceBERT Citation Networks Few-shot Node, Link Link
ARC[liu2024ARC] GCN Pre-training Conditional High - Citation/Social Networks, etc. Few-shot Graph Anomaly Detection link
UniGraph[he2024unigraph] GAT Pre-training Conditional High Llama, vicuna-7B, etc. Citation/Molecular Networks, etc. Zero/Few-shot Node, Edge, Graph Link
TEA-GLM[wang2024TeaGLM] GraphSAGE Pre-training Conditional High Vicuna-7B Citation/E-commerce Networks Zero-shot Node, Link link
GOFA[kong2025gofa] TransConv Pre-training Conditional High Llama2, Mistral, etc. Citation/Recommendation Networks Zero-shot Node, Link Link
HiGPT[tang2024higpt] HGNN Pre-training Conditional High Sentence-BERT Academic Networks Few-shot/Zero-shot Node Link
MDGPT[yu2024MDGPT] GIN Pre-training Open High - Citation/E-commerce Networks, etc. Few-shot Node, Graph -
AnyGraph[xia2024anygraph] GCN Pre-training Open High - E-commerce/Academic Networks, etc. Zero-shot Node, Link, Graph Link
GIMLET[zhao2023gimlet] - Pre-training Limited High T5 Molecular Networks Few-shot/Zero-shot Graph Link
GraphTranslator[zhang2024graphtranslator] GraphSAGE Pre-training Conditional High ChatGLM2-6B E-commerce/Academic Networks Few-shot/Zero-shot Node Link
GraphGPT[tang2024graphgpt] Graph Transformer Instruction tuning Conditional High Vicuna-7B-v1.1/1.5 Citation Networks Zero-shot Node, Link Link
GITA[wei2024gita] GCN, GAT Instruction tuning Open High GPT-4V, LLaVA-7B/13B Pure Structure Graphs Zero-shot/Full-shot Node, Link Link
InstructGraph[wang2024instructgraph] - Instruction tuning Conditional High LLaMA2 Pure Structure Graphs, Citation Networks, etc. Full/Few-shot Node, Generation, Reasoning, etc. Link
GraphCLIP[zhu2024graphclip] GCN, GAT Pre-traning Conditional High Qwen2-72B Academic/E-commerce Networks,etc Few/Zero-shot Node, Link, Graph Link
GraphWiz[chen2024graphwiz] - Instruction tuning Limited High LLaMA2-7B/13B, Mistral-7B Pure Structure Graphs Zero-shot/Few-shot Reasoning Link
GraphLoRA[yang2024graphlora] GAT PEFT Open High - E-commerce/Social Networks, etc. Full/Few-shot Node Link
OMOG[liu2024OMOG] SGC Pre-training Conditional High Sentence-BERT Citation/Social Networks, etc. Few/Zero-shot Node, Link -
MDP-GNN[lin2025MDPGNN] GT Pre-training Open Low - Citation/Social Networks, etc. Few-shot Node -
MDGFM[wang2025MDGFM] GCN Pre-training Open High - Citation/Social Networks, etc. Few-shot Node -
SAMGPT[yu2025SAMGPT] GCN Pre-training Open High - Citation/Social Networks, etc. Few-shot Node,Graph Link
UniGraph2[he2025unigraph2] GAT Pre-training Open High - Citation Networks, Knowledge Graphs, etc. Full/few-shot Node,Graph Link
GIT[wang2025GIT] GraphSAGE Pre-training Conditional Hard - Citation/Molecular Networks, etc. Full/few-shot Node, Link, Graph Link
MLDGG[tian2025mldgg] GCN Pre-training Open High - Social Networks Full-shot Node -
CMPGNN[wang2025bridgingCMPGNN] GCN, GIN Domain Adaptation Limited High - Citation/Social Networks Semi-supervised Node -
AEGOT-CDKT[wu2025crossAEGOTCDKT] GAT Domain Adaptation Limited Low - Educational Networks Supervised Node -
@@
TABLE I: A summary of models that leverage cross-domain knowledge to enhance graph-related tasks in the literature, ordered by their release time. Paradigm denotes the core technique used to achieve Cross-Domain Graph Learning. Domain indicates the domains across which knowledge is transferred. Scale and Difficulty refer to the cross-domain scales and the level of difficulty. Fine-tuning denotes whether it is necessary to fine-tune the parameters of Graph Neural Networks (GNNs). Prompting indicates the use of graph- or text-formatted prompts. Acronyms in Task: Node, Link, and Graph refer to node-level, link-level, and graph-level tasks, respectively. Beyond these general graph learning tasks, reasoning represents a higher-level reasoning ability on graphs, whereas other values correspond to specific application-oriented settings (e.g., graph anomaly detection).
VII Future Directions